### v1

# MoCrop: Training Free Motion Guided Cropping for Efficient Video Action Recognition

Abstract
We introduce MoCrop, a motion-aware adaptive cropping module for efficient video action recognition in the compressed domain. MoCrop uses motion vectors that are available in H.264 video to locate motion-dense regions and produces a single clip-level crop that is applied to all I-frames at inference. The module is training free, adds no parameters, and can be plugged into diverse backbones. A lightweight pipeline that includes denoising & merge (DM), Monte Carlo sampling (MCS), and adaptive cropping (AC) via a motion-density submatrix search yields robust crops with negligible overhead. On UCF101, MoCrop improves accuracy or reduces compute. With ResNet-50, it delivers +3.5% Top-1 accuracy at equal FLOPs (attention setting), or +2.4% Top-1 accuracy with 26.5% fewer FLOPs (efficiency setting). Applied to CoViAR, it reaches 89.2% Top-1 accuracy at the original cost and 88.5% Top-1 accuracy while reducing compute from 11.6 to 8.5 GFLOPs. Consistent gains on MobileNet-V3, EfficientNet-B1, and Swin-B indicate strong generality and make MoCrop practical for real-time deployment in the compressed domain. Our code and models are available at https://github.com/microa/MoCrop.
Index Terms—  Motion-Aware Cropping, Video Action Recognition, Motion Vectors, Computational Efficiency, Plug-and-Play Module
1 Introduction
Mirroring the human ability to rapidly ignore distractors and attend to informative regions, modern vision systems use attention to highlight critical image areas [1].
Deep learning transformed video action recognition: from hand-engineered features to learned spatiotemporal representations [2, 3]. Two-Stream underscored motion via precomputed optical flow [4], while C3D and video transformers learn end to end features and reach strong accuracy [5, 6, 7]. The cost is substantial computation, motivating more efficient approaches.
A complementary route is to exploit the compressed domain [8]: codecs such as H.264 store I-frames and represent others using motion vectors (MV) and residuals [9, 10]. These cues emphasize dynamic regions that often coincide with actions [11]. Systems like CoViAR process I-frames, motion vectors, and residuals in separate streams [8], and other works refine motion vectors toward optical flow [12]. In effect, compression already performs a form of feature selection [13].
Motivated by this, we hypothesize that a simple, parameter-free cropping strategy can capture spatially salient regions (SSR) using motion vectors alone.
We introduce MoCrop, a lightweight, training-free module that selects key spatial regions from motion vectors. We apply Monte Carlo downsampling [14], accumulate sampled motion vectors on a spatial grid, and derive a single crop for the clip. MoCrop attaches at inference, requires no model changes or retraining, and contrasts with many attention mechanisms that are embedded and trained end to end [15, 7]. Empirically, it improves accuracy while lowering inference FLOPs.
Scope and assumptions. MoCrop assumes access to codec motion vectors and focuses on a single crop per clip for temporal stability. The design is orthogonal to backbone choice, training protocol, and test-time augmentation, and it can be enabled or disabled without retraining.
Contributions.
  • A training-free, plug-and-play MV-driven adaptive cropping module.
  • Accuracy and efficiency gains across common backbones.
  • A lightweight design using Monte Carlo downsampling on motion vectors to reduce computation while capturing discriminative regions.
2 Method
Fig. 1: MoCrop pipeline. Motion vectors identify spatially salient regions (SSR) to guide I-frame cropping; cropped frames are fed to the action recognition model.
MoCrop is a preprocessing module that uses motion vectors from compressed streams to locate spatially salient regions (SSR) and compute a single adaptive crop, focusing inference on informative regions. This reduces input resolution and FLOPs and often improves accuracy by suppressing background distraction.
2.1 Problem Formulation
Given decoded I-frames and raw motion vectors, let $V_{\rm raw}=\{(p_{k},\Delta p_{k})\}_{k=1}^{T}$ with $p_{k}=(x_{k},y_{k})$ the origin and $\Delta p_{k}$ the displacement. We seek a function $f$ that maps motion vectors to a clip-level box $B^{\ast}$:
$B^{\ast}=f(V_{\rm raw}),$
computed once per clip for temporal consistency. Each I-frame $I_{i}$ is then cropped by
$I_{i}^{\prime}=g(B^{\ast},I_{i}).$
2.2 Detailed Algorithm
MoCrop comprises three processing modules, DM (denoise & merge), MCS (Monte Carlo sampling), and AC (adaptive cropping). DM suppresses static/low-amplitude MVs and merges them across frames; MCS randomly subsamples the filtered MV to a fixed budget and counts them per cell to form the motion-density map (MD map); AC aggregates MD on a grid, searches a high-score rectangle near a target area ratio, and crops the clip accordingly.
2.2.1 Stage 1: Denoise-and-Merge (DM)
Extraction and Filtering. Let $V_{\rm raw}=\{(p_{k},\Delta p_{k})\}_{k=1}^{T}$. We suppress static/near-static vectors:
$V_{\rm filtered}=\{(p,\Delta p)\in V_{\rm raw}:\ \|\Delta p\|_{2}>\epsilon\},$
with small threshold $\epsilon$.
2.2.2 Stage 2: Monte Carlo Sampling & MD Map
Monte Carlo Sampling (MCS). For efficiency when $|V_{\rm filtered}|$ is large, we uniformly sample a subset $V_{\rm sampled}\subseteq V_{\rm filtered}$ of size $|V_{\rm sampled}|=\min(N,|V_{\rm filtered}|)$, preserving spatial distribution while reducing points.
MD Map. We discretize a frame of size $(W_{\rm video},H_{\rm video})$ into an $h\times w$ grid. Cell $C_{ij}$ spans
$C_{ij}=\Big[j\tfrac{W_{\rm video}}{w},(j+1)\tfrac{W_{\rm video}}{w}\Big)\times\Big[i\tfrac{H_{\rm video}}{h},(i+1)\tfrac{H_{\rm video}}{h}\Big).$
The map $M$ counts sampled motion vectors per cell:
$M_{ij}=\sum_{(p_{k},\cdot)\in V_{\rm sampled}}I(p_{k}\in C_{ij}),$
where $I(\cdot)$ is the indicator; larger $M_{ij}$ indicates stronger motion.
2.2.3 Stage 3: Submatrix Search and Box Generation (AC)
We find the submatrix with maximal motion mass:
$(i^{\ast},j^{\ast},H_{s}^{\ast},W_{s}^{\ast})=\mathop{\rm arg\,max}_{i,j,H_{s},W_{s}}\ \sum_{p=i}^{i+H_{s}-1}\ \sum_{q=j}^{j+W_{s}-1}M_{pq}$
subject to
$\begin{array}[]{l}1\leq H_{s}\leq h,\quad 1\leq W_{s}\leq w,\\[2.0pt] 0\leq i\leq h-H_{s},\quad 0\leq j\leq w-W_{s},\\[2.0pt] (1-\delta)\alpha hw\leq H_{s}W_{s}\leq(1+\delta)\alpha hw,\end{array}$
with target area ratio $\alpha\in(0,1]$ and tolerance $\delta$ (e.g., $0.1$). We convert grid parameters to a normalized YOLO [16] box $(c_{x},c_{y},w_{b},h_{b})$:
$c_{x}=\frac{j^{\ast}+W_{s}^{\ast}/2}{w},\ c_{y}=\frac{i^{\ast}+H_{s}^{\ast}/2}{h},\ w_{b}=\frac{W_{s}^{\ast}}{w},\ h_{b}=\frac{H_{s}^{\ast}}{h}.$
Here $i,j,H_{s},W_{s}\in\mathbb{N}$ denote integer grid indices and sizes.
2.2.4 Stage 4: Cropping I-frames
We map the normalized box to pixel coordinates $(x_{1},y_{1},x_{2},y_{2})$:
$\begin{array}[]{l}x_{1}=\max\Big(0,\ \lfloor(c_{x}-w_{b}/2)\,W_{\rm video}\rfloor\Big),\\[3.0pt] y_{1}=\max\Big(0,\ \lfloor(c_{y}-h_{b}/2)\,H_{\rm video}\rfloor\Big),\\[3.0pt] x_{2}=\min\Big(W_{\rm video},\ \lfloor(c_{x}+w_{b}/2)\,W_{\rm video}\rfloor\Big),\\[3.0pt] y_{2}=\min\Big(H_{\rm video},\ \lfloor(c_{y}+h_{b}/2)\,H_{\rm video}\rfloor\Big).\end{array}$
and extract
$I_{\rm crop}=I[y_{1}:y_{2},\ x_{1}:x_{2}].$
The same box is applied to all I-frames in the clip, yielding an efficient, training-free, plug-and-play preprocessor.
Components. DM: denoise & temporal merge of motion vectors; MCS: Monte Carlo sampling on filtered MVs to estimate a motion-density (MD) map on an $h{\times}w$ grid; AC: grid-wise aggregation and rectangle search over MD to produce a single clip-level crop (Fig. 1).
2.3 Complexity and Practicality
Complexity. DM scans raw motion vectors once to threshold and (optionally) merge across frames, i.e., $O(|V_{\rm raw}|)$ time.
Building the motion-density map costs $O(|V_{\rm sampled}|+hw)$ for counting on an $h\times w$ grid. For submatrix search, our prototype directly sums each candidate window on an $R{\times}C$ grid (R=hR{=}h, C=wC{=}w): the worst-case time is $O(R^{3}C^{3})$ (looping over $(H_{s},W_{s})$ and positions with $O(H_{s}W_{s})$ sum per window). In practice $R{\times}C\!\leq\!48$ (e.g., $6{\times}8$), so this overhead is negligible compared with backbone inference. Optionally, a 2D integral image reduces each window score to $O(1)$, yielding $O(\#\text{cands})$ overall; when $(H_{s},W_{s})$ are fixed, a sliding-window implementation is $O(hw)$ total.
Hyperparameters. $\epsilon$ can be set by a small absolute magnitude or by a low percentile of $\|\Delta p\|_{2}$; $N$ trades variance and cost and is robust across a range; $(h,w)$ can be chosen proportional to frame size and codec macroblocks; $(\alpha,\delta)$ control the crop area and tolerance. For strong camera motion, subtracting a global estimate (for example, the median motion vector) stabilizes the map. When the map is flat, falling back to a center crop is a safe default.
3 Experiments
3.1 Dataset and Metrics
All experiments are conducted on UCF101, a widely used benchmark for human action recognition featuring 13,320 videos across 101 action classes [17]. Given that our approach is designed for efficiency, we evaluate performance by jointly considering Top-1 accuracy and computational cost in GFLOPs.
Table 1: Demonstration that MoCrop provides dual advantages as a plug-and-play module on modern backbones using the UCF101 dataset. The table highlights the performance gains in both the accuracy-focused Attention Setting and the cost-saving Efficiency Setting.
Backbone Dataset Variant Top-1 Acc. (%) GFLOPs $\Delta$ Acc. (%) $\Delta$ FLOPs (%)
ResNet-50 [18] UCF101 (A) Baseline (224px) 80.1 4.11 - -
(B) + MoCrop (Efficiency, 192px) 82.5 3.02 +2.4 -26.5
(C) + MoCrop (Attention, 224px) 83.6 4.11 +3.5 0.0
MobileNet-V3 [19] UCF101 (A) Baseline (224px) 78.3 0.22 - -
(B) + MoCrop (Efficiency, 192px) 79.3 0.17 +1.0 -22.7
(C) + MoCrop (Attention, 224px) 80.2 0.22 +1.9 0.0
EfficientNet-B1 [20] UCF101 (A) Baseline (224px) 82.1 0.59 - -
(B) + MoCrop (Efficiency, 192px) 83.6 0.43 +1.5 -27.1
(C) + MoCrop (Attention, 224px) 84.5 0.59 +2.4 0.0
Swin-B [6] UCF101 (A) Baseline (224px) 87.3 15.5 - -
(B) + MoCrop (Efficiency, 192px) 87.2 12.6 -0.1 -18.7
(C) + MoCrop (Attention, 224px) 87.5 15.5 +0.2 0.0
3.2 Experimental Setup
We evaluate MoCrop on the official UCF101 split. Following standard practice [8], we sample three frames per video for training and use typical data augmentation. At inference, we freeze the model weights and evaluate two main settings: Baseline: The model trained on UCF101 is evaluated on the standard test set with a 224px input. MoCrop: The same model is evaluated after each video is pre-processed by MoCrop. We test two variants:
  • Efficiency Setting: MoCrop crops the video, which is then resized to a smaller resolution (192px) to reduce FLOPs while aiming to maintain or improve accuracy.
  • Attention Setting: MoCrop crops the video, which is resized back to the original resolution (224px). This isolates the accuracy improvement gained by focusing on salient regions, without changing the computational cost.
3.3 Performance and Efficiency Analysis
Main Results.
Table 1 demonstrates the powerful dual advantages of MoCrop as a plug-and-play module for various modern backbones. In the Attention Setting, where FLOPs are held constant, MoCrop consistently boosts performance by forcing the model to concentrate on motion-rich areas. The most significant gain is seen with ResNet-50 [18], which improves by a remarkable +3.5% Top-1 accuracy. Other models like EfficientNet-B1 [20] (+2.4%) and MobileNet-V3 [19] (+1.9%) also show substantial improvements.
In the Efficiency Setting, MoCrop proves its ability to reduce computational costs and often enhances accuracy. For instance, ResNet-50 sees a +2.4% accuracy increase while FLOPs are cut by a massive 26.5%. Even for the heavy Swin-B Transformer [6], where accuracy gains are harder to achieve, our method reduces the computational load by 18.7% with only a negligible 0.1% accuracy dip. These results confirm that MoCrop offers an exceptional and highly flexible trade-off between accuracy and efficiency.
3.4 Ablation Studies
To validate the contributions of the internal components of MoCrop and its effectiveness as a general module, we present a series of ablation studies in Table 2. The studies are organized into three parts: (a) an incremental analysis of our module’s components, (b) a comparison against standard cropping methods, and (c) an application of MoCrop to enhance a prior state-of-the-art method.
Incremental Component Analysis. Using ResNet-18 [18] as a base, we find that each component of MoCrop contributes to the final performance. Simply adding Adaptive Cropping (AC) improves accuracy from 79.3% to 80.6%. The addition of Denoise & Merge (DM) and Monte Carlo Sampling (MCS; MD map) further pushes the accuracy to a final 81.4%. This confirms that all parts of our design are integral to its success.
Comparison with Cropping Strategies. MoCrop significantly outperforms standard cropping techniques. While 90% Random and Center crops offer a slight improvement over the baseline, our adaptive method achieves a much higher accuracy of 81.4% with fewer FLOPs. This demonstrates the superiority of using motion information to guide cropping decisions over static, fixed-ratio approaches.
Enhancing Prior Works. To show its versatility, we applied MoCrop to CoViAR, a well-known compressed-domain method. The results are compelling. In the efficiency setting, MoCrop improves CoViAR’s accuracy to 88.5% while reducing its FLOPs from 11.6 GFLOPs to just 8.5 GFLOPs. In the attention setting, it boosts the accuracy to an impressive 89.2%. This shows that MoCrop is not just a standalone module but can also serve as a powerful enhancer for other methods in the field.
Table 2: Ablation studies and comparative analysis on ResNet-18 and CoViAR. MoCrop shows clear benefits over baselines and standard techniques, and effectively enhances prior state-of-the-art methods.
Setup Method Acc. (%) GFLOPs
(a) Ablation on MoCrop Components (on ResNet-18 [18])
ResNet-18 [18] Baseline 79.3 1.82
+ AC 80.6 1.54
+ AC + DM 81.3 1.54
+ AC + DM + MCS (Full) 81.4 1.54
(b) Comparison with Cropping Strategies (on ResNet-18)
ResNet-18 Baseline 79.3 1.82
Random Crop (90%) 80.4 1.73
Center Crop (90%) 80.2 1.73
MoCrop (Adaptive) 81.4 1.54
(c) Enhancing Prior Works (on UCF101, I-frames only)
CoViAR [8] Baseline (Reported in [8]) 87.7 11.6
+ MoCrop (Efficiency, 192px) 88.5 8.5
+ MoCrop (Attention, 224px) 89.2 11.6
(a) PullUps
(b) ApplyEyeMakeup
(c) JumpRope
(d) Archery
(e) BalanceBeam
(f) HighJump
Fig. 2: Visualization of MoCrop (green box) vs. center cropping (red box), alongside motion-density (MD) maps. Each subfigure presents two views in order: RGB frame with boxes (left) and the MD map (right), shown across six UCF101 classes.
3.5 Visualization
Fig. 2 presents qualitative examples of adaptive cropping with MoCrop. For each class, the left panel shows an RGB frame with the green box predicted by MoCrop and the red fixed center-crop box; the right panel shows the corresponding motion-density (MD) map. Across six UCF101 classes (PullUps, ApplyEyeMakeup, JumpRope, Archery, BalanceBeam, and HighJump), MoCrop places a tighter crop around the actor and the regions with salient motion (e.g., the torso during pull-ups or the take-off area in high jump). In contrast, the center crop often includes large static background or truncates the subject (notably in Archery and BalanceBeam). The MD maps on the right concentrate energy near the actor and align with the green crop, illustrating the signal that guides our policy.
4 Discussion and Conclusion
We present MoCrop, a lightweight, training-free module using compressed-domain motion vectors to generate a single, clip-level crop of motion-dense regions. By focusing inference, MoCrop reduces resolution and FLOPs, often boosting accuracy, and integrates with existing backbones without retraining.
When it helps. MoCrop is most effective when actions coincide with localized motion and when background clutter dilutes fixed crops. The map-based formulation is robust to moderate noise after filtering and sampling, and it generalizes across backbones since it operates as a preprocessing step.
Limitations. MoCrop relies on codec-provided motion vectors (e.g., H.264) and is less informative for raw videos unless motion estimation is run in parallel. This dependency is acceptable since most datasets and streaming pipelines use compressed video. Large camera motion or scene-wide dynamics can bias the density map; mitigations include global-motion suppression, temporal smoothing, and combining motion vector magnitude with residual cues.
Future work. We plan to extend MoCrop to diagonally elongated or multi-region actions via multi-box proposals, explore codec-agnostic motion signals, and validate on larger datasets and backbones. We will also study learned selection over MoCrop proposals while keeping inference cost low.
In summary, MoCrop offers a practical plug-and-play path to efficiency and accuracy in compressed-domain video understanding, requiring no additional parameters and minimal compute overhead.

### v2

@@
 Abstract
-We introduce MoCrop, a motion-aware adaptive cropping module for efficient video action recognition in the compressed domain.
-MoCrop uses motion vectors that are available in H.264 video to locate motion-dense regions and produces a single clip-level crop that is applied to all I-frames at inference.
-The module is training free, adds no parameters, and can be plugged into diverse backbones.
-A lightweight pipeline that includes denoising & merge (DM), Monte Carlo sampling (MCS), and adaptive cropping (AC) via a motion-density submatrix search yields robust crops with negligible overhead.
-On UCF101, MoCrop improves accuracy or reduces compute.
-With ResNet-50, it delivers +3.5% Top-1 accuracy at equal FLOPs (attention setting), or +2.4% Top-1 accuracy with 26.5% fewer FLOPs (efficiency setting).
-Applied to CoViAR, it reaches 89.2% Top-1 accuracy at the original cost and 88.5% Top-1 accuracy while reducing compute from 11.6 to 8.5 GFLOPs.
-Consistent gains on MobileNet-V3, EfficientNet-B1, and Swin-B indicate strong generality and make MoCrop practical for real-time deployment in the compressed domain.
+Standard video action recognition models often process typically resized full frames, suffering from spatial redundancy and high computational costs.
+To address this, we introduce MoCrop, a motion-aware adaptive cropping module designed for efficient video action recognition in the compressed domain.
+Leveraging Motion Vectors (MVs) naturally available in H.264 video, MoCrop localizes motion-dense regions to produce adaptive crops at inference without requiring any training or parameter updates.
+Our lightweight pipeline synergizes three key components: Merge & Denoise (MD) for outlier filtering, Monte Carlo Sampling (MCS) for efficient importance sampling, and Motion Grid Search (MGS) for optimal region localization.
+This design allows MoCrop to serve as a versatile “plug-and-play” module for diverse backbones.
+Extensive experiments on UCF101 demonstrate that MoCrop serves as both an accelerator and an enhancer.
+With ResNet-50, it achieves a +3.5% boost in Top-1 accuracy at equivalent FLOPs (Attention Setting), or a +2.4% accuracy gain with 26.5% fewer FLOPs (Efficiency Setting).
+When applied to CoViAR, it improves accuracy to 89.2% or reduces computation by roughly 27% (from 11.6 to 8.5 GFLOPs).
+Consistent gains across MobileNet-V3, EfficientNet-B1, and Swin-B confirm its strong generality and suitability for real-time deployment.
 Our code and models are available at https://github.com/microa/MoCrop.
-Index Terms—  Motion-Aware Cropping, Video Action Recognition, Motion Vectors, Computational Efficiency, Plug-and-Play Module
-1 Introduction
-Mirroring the human ability to rapidly ignore distractors and attend to informative regions, modern vision systems use attention to highlight critical image areas [1].
-Deep learning transformed video action recognition: from hand-engineered features to learned spatiotemporal representations [2, 3].
-Two-Stream underscored motion via precomputed optical flow [4], while C3D and video transformers learn end to end features and reach strong accuracy [5, 6, 7].
-The cost is substantial computation, motivating more efficient approaches.
-A complementary route is to exploit the compressed domain [8]: codecs such as H.264 store I-frames and represent others using motion vectors (MV) and residuals [9, 10].
-These cues emphasize dynamic regions that often coincide with actions [11].
-Systems like CoViAR process I-frames, motion vectors, and residuals in separate streams [8], and other works refine motion vectors toward optical flow [12].
-In effect, compression already performs a form of feature selection [13].
-Motivated by this, we hypothesize that a simple, parameter-free cropping strategy can capture spatially salient regions (SSR) using motion vectors alone.
-We introduce MoCrop, a lightweight, training-free module that selects key spatial regions from motion vectors.
-We apply Monte Carlo downsampling [14], accumulate sampled motion vectors on a spatial grid, and derive a single crop for the clip.
-MoCrop attaches at inference, requires no model changes or retraining, and contrasts with many attention mechanisms that are embedded and trained end to end [15, 7].
-Empirically, it improves accuracy while lowering inference FLOPs.
-Scope and assumptions.
-MoCrop assumes access to codec motion vectors and focuses on a single crop per clip for temporal stability.
-The design is orthogonal to backbone choice, training protocol, and test-time augmentation, and it can be enabled or disabled without retraining.
-Contributions.
-  • A training-free, plug-and-play MV-driven adaptive cropping module.
-  • Accuracy and efficiency gains across common backbones.
-  • A lightweight design using Monte Carlo downsampling on motion vectors to reduce computation while capturing discriminative regions.
-2 Method
-Fig. 1: MoCrop pipeline. Motion vectors identify spatially salient regions (SSR) to guide I-frame cropping; cropped frames are fed to the action recognition model.
-MoCrop is a preprocessing module that uses motion vectors from compressed streams to locate spatially salient regions (SSR) and compute a single adaptive crop, focusing inference on informative regions.
+I Introduction
+Deep learning has fundamentally transformed video action recognition, evolving from hand-crafted features to sophisticated spatiotemporal representations learned via 3D convolutions and Video Transformers [12, 15, 8, 1].
+However, this performance leap comes with substantial computational costs.
+A primary source of inefficiency is spatial redundancy: standard models process the entire frame—often resized to a fixed resolution (e.g., $224\times 224$)—regardless of whether the action occurs in a localized region or spans the whole scene.
+While spatial attention mechanisms [3, 17] can mitigate this by highlighting informative areas, they typically require end-to-end training, introduce additional parameters, and are coupled tightly with specific backbone architectures.
+To address efficiency without sacrificing accuracy, the compressed domain offers a compelling alternative [18, 16].
+Video codecs like H.264 naturally encode motion information via Motion Vectors (MVs) to reduce temporal redundancy.
+Prior works such as CoViAR [18] and MV2Flow [6] treat these MVs as input modalities to learn features.
+In contrast, we hypothesize that MVs serve a more primal purpose: they are essentially ”free” saliency maps that can guide the model to focus on Spatially Salient Regions (SSR) before inference begins [20, 10].
+Building on this insight, we introduce MoCrop, a lightweight, training-free module that leverages codec MVs for adaptive spatial cropping.
+Unlike learned attention modules, MoCrop operates as a pre-processor.
+It addresses the inherent noisiness and sparsity of raw MVs through a novel pipeline: Merge & Denoise (MD) to filter outliers, Monte Carlo Sampling (MCS) to drastically reduce processing overhead, and Motion Grid Search (MGS) to lock onto the optimal action region.
+MoCrop offers a versatile ”Plug-and-Play” capability.
+In an Efficiency Setting, it allows models to process smaller, cropped resolutions (e.g., $192\text{px}$), significantly reducing FLOPs.
+In an Attention Setting, it removes background clutter for standard inputs, boosting accuracy.
+Crucially, MoCrop requires no retraining of the backbone model, making it immediately applicable to diverse architectures ranging from ResNet to Swin Transformers.
+Our main contributions are summarized as follows:
+  • We propose MoCrop, a training-free, parameter-free adaptive cropping module that utilizes compressed-domain Motion Vectors to guide efficient video recognition.
+  • We design a high-efficiency pipeline incorporating Monte Carlo Sampling and dual-objective grid search, enabling robust region selection with negligible computational overhead.
+  • We demonstrate that MoCrop serves as a universal accelerator and enhancer.
+Experiments on UCF101 show it can improve Top-1 accuracy by up to 3.5% (ResNet-50) or reduce FLOPs by over 26% while maintaining accuracy, generalizing across CNNs and Transformers.
+II Related Work
+Video Action Recognition.
+Early approaches to video action recognition relied on hand-crafted features but struggled with scalability [19].
+The emergence of deep learning marked a paradigm shift, with the Two-Stream Network [11] leveraging pre-computed optical flow to capture motion patterns.
+However, optical flow computation remains expensive.
+C3D [15] introduced 3D convolutions to learn spatiotemporal features end-to-end, avoiding explicit optical flow but incurring high computational costs.
+More recently, video transformers such as Video Swin [8] and TimeSformer [1] achieve state-of-the-art performance by modeling long-range dependencies through self-attention, though at significant computational expense.
+These methods motivate efficient alternatives that reduce computation while maintaining accuracy.
+Compressed-Domain Video Understanding.
+Compressed video formats like H.264 encode videos using I-frames (independently coded), P-frames (predicted from previous frames), and B-frames (bi-directionally predicted), with MVs and residuals encoding inter-frame differences [16].
+Leveraging these codec-provided features offers computational advantages over processing raw pixels.
+CoViAR [18] pioneered compressed-domain action recognition by processing I-frames, MVs, and residuals in separate streams.
+MV2Flow [6] refines MVs to approximate optical flow, bridging efficiency and accuracy.
+Recent works [9, 10] further exploit MVs to guide temporal frame selection and spatial attention, demonstrating that compressed-domain cues can enhance both efficiency and recognition performance.
+Spatial Attention and Adaptive Cropping.
+Spatial attention mechanisms identify and emphasize informative regions within frames.
+Early convolutional approaches like CBAM [17] and SE-Net [7] apply channel and spatial attention modules trained end-to-end.
+Transformer-based methods [1, 2] use self-attention to model spatial dependencies but require substantial computation and parameter overhead.
+In the context of video understanding, adaptive cropping has been explored to focus on action-relevant regions.
+Traditional methods include center cropping, random cropping during training, and saliency-based ROI extraction [3].
+Recent work by Shahabinejad et al.
+[10] integrates MVs into learned spatial attention for compressed video, achieving improved efficiency.
+However, most attention and cropping methods require training or fine-tuning for each backbone and dataset.
+Figure 1: MoCrop pipeline. MVs identify SSR to guide I-frame cropping; cropped frames are fed to the action recognition model.
+Our Approach.
+Unlike prior attention mechanisms that are learned end-to-end or require retraining, MoCrop is a training-free, plug-and-play module that operates solely at inference time.
+By aggregating MVs into a motion-density map and performing a lightweight spatial search, MoCrop adaptively crops to motion-salient regions without modifying model parameters or training protocols.
+This design makes it compatible with diverse backbones and datasets, offering a practical path to efficiency and accuracy improvements in compressed-domain video action recognition.
+III Method
+MoCrop is a preprocessing module that uses MVs from compressed streams to locate SSR and compute adaptive crops, focusing inference on informative regions.
 This reduces input resolution and FLOPs and often improves accuracy by suppressing background distraction.
-2.1 Problem Formulation
-Given decoded I-frames and raw motion vectors, let $V_{\rm raw}=\{(p_{k},\Delta p_{k})\}_{k=1}^{T}$ with $p_{k}=(x_{k},y_{k})$ the origin and $\Delta p_{k}$ the displacement.
-We seek a function $f$ that maps motion vectors to a clip-level box $B^{\ast}$:
+III-A Overview
+Fig.
+1 illustrates the complete MoCrop pipeline.
+MoCrop comprises three processing modules: Merge & Denoise (MD), Monte Carlo Sampling (MCS), and Motion Grid Search (MGS).
+MD retains top-$k$ MVs by magnitude; MCS applies weighted importance sampling to downsample filtered MVs and builds a motion-density map; MGS searches for a high-score rectangle via dual-objective optimization to determine the cropping region.
+III-B Problem Formulation
+Given decoded I-frames and raw MVs, let $V_{\rm raw}=\{(p_{k},\Delta p_{k})\}_{k=1}^{T}$ with $p_{k}=(x_{k},y_{k})$ the origin and $\Delta p_{k}$ the displacement.
+We seek a function $f$ that maps MVs to a bounding box $B^{\ast}$:
 $B^{\ast}=f(V_{\rm raw}),$
-computed once per clip for temporal consistency. Each I-frame $I_{i}$ is then cropped by
+and each I-frame $I_{i}$ is then cropped by
 $I_{i}^{\prime}=g(B^{\ast},I_{i}).$
-2.2 Detailed Algorithm
-MoCrop comprises three processing modules, DM (denoise & merge), MCS (Monte Carlo sampling), and AC (adaptive cropping).
-DM suppresses static/low-amplitude MVs and merges them across frames; MCS randomly subsamples the filtered MV to a fixed budget and counts them per cell to form the motion-density map (MD map); AC aggregates MD on a grid, searches a high-score rectangle near a target area ratio, and crops the clip accordingly.
-2.2.1 Stage 1: Denoise-and-Merge (DM)
-Extraction and Filtering. Let $V_{\rm raw}=\{(p_{k},\Delta p_{k})\}_{k=1}^{T}$. We suppress static/near-static vectors:
-$V_{\rm filtered}=\{(p,\Delta p)\in V_{\rm raw}:\ \|\Delta p\|_{2}>\epsilon\},$
-with small threshold $\epsilon$.
-2.2.2 Stage 2: Monte Carlo Sampling & MD Map
-Monte Carlo Sampling (MCS).
-For efficiency when $|V_{\rm filtered}|$ is large, we uniformly sample a subset $V_{\rm sampled}\subseteq V_{\rm filtered}$ of size $|V_{\rm sampled}|=\min(N,|V_{\rm filtered}|)$, preserving spatial distribution while reducing points.
-MD Map. We discretize a frame of size $(W_{\rm video},H_{\rm video})$ into an $h\times w$ grid. Cell $C_{ij}$ spans
-$C_{ij}=\Big[j\tfrac{W_{\rm video}}{w},(j+1)\tfrac{W_{\rm video}}{w}\Big)\times\Big[i\tfrac{H_{\rm video}}{h},(i+1)\tfrac{H_{\rm video}}{h}\Big).$
-The map $M$ counts sampled motion vectors per cell:
-$M_{ij}=\sum_{(p_{k},\cdot)\in V_{\rm sampled}}I(p_{k}\in C_{ij}),$
-where $I(\cdot)$ is the indicator; larger $M_{ij}$ indicates stronger motion.
-2.2.3 Stage 3: Submatrix Search and Box Generation (AC)
-We find the submatrix with maximal motion mass:
-$(i^{\ast},j^{\ast},H_{s}^{\ast},W_{s}^{\ast})=\mathop{\rm arg\,max}_{i,j,H_{s},W_{s}}\ \sum_{p=i}^{i+H_{s}-1}\ \sum_{q=j}^{j+W_{s}-1}M_{pq}$
-subject to
-$\begin{array}[]{l}1\leq H_{s}\leq h,\quad 1\leq W_{s}\leq w,\\[2.0pt] 0\leq i\leq h-H_{s},\quad 0\leq j\leq w-W_{s},\\[2.0pt] (1-\delta)\alpha hw\leq H_{s}W_{s}\leq(1+\delta)\alpha hw,\end{array}$
-with target area ratio $\alpha\in(0,1]$ and tolerance $\delta$ (e.g., $0.1$). We convert grid parameters to a normalized YOLO [16] box $(c_{x},c_{y},w_{b},h_{b})$:
-$c_{x}=\frac{j^{\ast}+W_{s}^{\ast}/2}{w},\ c_{y}=\frac{i^{\ast}+H_{s}^{\ast}/2}{h},\ w_{b}=\frac{W_{s}^{\ast}}{w},\ h_{b}=\frac{H_{s}^{\ast}}{h}.$
-Here $i,j,H_{s},W_{s}\in\mathbb{N}$ denote integer grid indices and sizes.
-2.2.4 Stage 4: Cropping I-frames
-We map the normalized box to pixel coordinates $(x_{1},y_{1},x_{2},y_{2})$:
-$\begin{array}[]{l}x_{1}=\max\Big(0,\ \lfloor(c_{x}-w_{b}/2)\,W_{\rm video}\rfloor\Big),\\[3.0pt] y_{1}=\max\Big(0,\ \lfloor(c_{y}-h_{b}/2)\,H_{\rm video}\rfloor\Big),\\[3.0pt] x_{2}=\min\Big(W_{\rm video},\ \lfloor(c_{x}+w_{b}/2)\,W_{\rm video}\rfloor\Big),\\[3.0pt] y_{2}=\min\Big(H_{\rm video},\ \lfloor(c_{y}+h_{b}/2)\,H_{\rm video}\rfloor\Big).\end{array}$
-and extract
-$I_{\rm crop}=I[y_{1}:y_{2},\ x_{1}:x_{2}].$
+III-C Detailed Algorithm
+We now describe each stage in detail.
+III-C1 Stage 1: Merge & Denoise (MD)
+Extraction and Filtering. Let $V_{\rm raw}=\{(p_{k},\Delta p_{k})\}_{k=1}^{T}$ denote raw MVs across frames. We retain only the top-$k$ vectors by motion magnitude:
+$V_{\rm filtered}=\text{top-}k(V_{\rm raw},\alpha),\quad k=\lfloor\alpha|V_{\rm raw}|\rfloor,$
+where $\alpha\in(0,1]$ controls the denoising strength (e.g., $\alpha{=}0.01$ keeps the top 1%).
+III-C2 Stage 2: Monte Carlo Sampling (MCS)
+For efficiency when $|V_{\text{filtered}}|$ is large, we apply weighted importance sampling. Each vector is sampled with probability proportional to its motion magnitude raised to power $\beta$:
+$p_{i}$ $\propto\|\Delta p_{i}\|_{2}^{\beta},$
+$V_{\text{sampled}}$ $\sim\text{Multinomial}(V_{\text{filtered}},N,\{p_{i}\}),$
+where $N{=}\lfloor\gamma|V_{\text{filtered}}|\rfloor$, $\gamma{\in}(0,1]$ is the sampling ratio (e.g., $\gamma{=}0.1$ samples 10%), and $\beta{>}0$ (e.g., $\beta{=}4.0$) biases toward high-magnitude motion.
+III-C3 Stage 3: Motion Grid Search (MGS)
+Grid Quantization. We discretize the frame into an $h\times w$ grid and map each MV to its corresponding cell:
+$M_{ij}=\sum_{(p_{k},\cdot)\in V_{\text{sampled}}}\mathbb{I}(p_{k}\in C_{ij}),$
+where $C_{ij}$ spans $j\frac{W}{w},(j+1)\frac{W}{w})\times[i\frac{H}{h},(i+1)\frac{H}{h})$. This motion-density map $M$ guides the subsequent search.
+Submatrix Search. We identify the grid region $R$ with maximum weighted score:
+$R^{*}$ $=\arg\max_{R\in\mathcal{R}_{\rho}}S(R),$
+$S(R)$ $=w_{\text{sum}}\cdot\sum_{(i,j)\in R}M_{ij}+w_{\text{avg}}\cdot\frac{1}{|R|}\sum_{(i,j)\in R}M_{ij},$
+where $\mathcal{R}_{\rho}{=}\{R:(1{-}\delta)\rho hw{\leq}|R|{\leq}(1{+}\delta)\rho hw\}$ constrains candidate rectangles to target area ratio $\rho$ with tolerance $\delta$, and $(w_{\text{sum}},w_{\text{avg}})$ balance total vs.
+average density (e.g., $(0.6,0.4)$).
+The optimal region $R^{*}$ is converted to normalized box coordinates for cropping.
+III-C4 Stage 4: Cropping I-frames
+We map the grid region $R^{*}$ to pixel coordinates and extract the corresponding crop:
+$I_{\rm crop}$ $=I[y_{1}:y_{2},\ x_{1}:x_{2}],$
+$(x_{1},y_{1},x_{2},y_{2})$ $=\text{Grid2Pixel}(R^{*},W_{\rm video},H_{\rm video}).$
 The same box is applied to all I-frames in the clip, yielding an efficient, training-free, plug-and-play preprocessor.
-Components.
-DM: denoise & temporal merge of motion vectors; MCS: Monte Carlo sampling on filtered MVs to estimate a motion-density (MD) map on an $h{\times}w$ grid; AC: grid-wise aggregation and rectangle search over MD to produce a single clip-level crop (Fig.
-1).
-2.3 Complexity and Practicality
-Complexity. DM scans raw motion vectors once to threshold and (optionally) merge across frames, i.e., $O(|V_{\rm raw}|)$ time.
+III-D Complexity and Practicality
+MD performs top-$k$ selection via argpartition in $O(|V_{\rm raw}|)$ time.
+MCS computes magnitude weights and samples in $O(|V_{\rm filtered}|)$.
 Building the motion-density map costs $O(|V_{\rm sampled}|{+}hw)$ for counting on an $h{\times}w$ grid.
-For submatrix search, our prototype directly sums each candidate window on an $R{\times}C$ grid (R=hR{=}h, C=wC{=}w): the worst-case time is $O(R^{3}C^{3})$ (looping over $(H_{s},W_{s})$ and positions with $O(H_{s}W_{s})$ sum per window).
-In practice $R{\times}C\!\leq\!48$ (e.g., $6{\times}8$), so this overhead is negligible compared with backbone inference.
-Optionally, a 2D integral image reduces each window score to $O(1)$, yielding $O(\#\text{cands})$ overall; when $(H_{s},W_{s})$ are fixed, a sliding-window implementation is $O(hw)$ total.
+MGS enumerates all rectangles within the area constraint: with integral images, scoring each candidate is $O(1)$, and the total search is $O(h^{2}w^{2})$ but negligible for small grids (e.g., $16{\times}9$).
+Overall preprocessing overhead is negligible, making MoCrop practical for real-time deployment.
 Hyperparameters.
-$\epsilon$ can be set by a small absolute magnitude or by a low percentile of $\|\Delta p\|_{2}$; $N$ trades variance and cost and is robust across a range; $(h,w)$ can be chosen proportional to frame size and codec macroblocks; $(\alpha,\delta)$ control the crop area and tolerance.
-For strong camera motion, subtracting a global estimate (for example, the median motion vector) stabilizes the map.
-When the map is flat, falling back to a center crop is a safe default.
-3 Experiments
-3.1 Dataset and Metrics
+Key parameters include: $\alpha$ (denoising ratio, e.g., 0.01) controls noise filtering strength; $\gamma$ (sampling ratio, e.g., 0.1) trades variance for speed; $\beta$ (importance weight, e.g., 4.0) biases sampling toward high-motion regions; $(h,w)$ (grid resolution, e.g., $16{\times}9$) balances localization precision and search cost; $\rho$ (target area ratio) sets crop size; $(w_{\text{sum}},w_{\text{avg}})$ (e.g., 0.6, 0.4) balance region compactness vs.
+density.
+IV Experiments
+IV-A Dataset and Metrics
 All experiments are conducted on UCF101, a widely used benchmark for human action recognition featuring 13,320 videos across 101 action classes [13].
 Given that our approach is designed for efficiency, we evaluate performance by jointly considering Top-1 accuracy and computational cost in GFLOPs.
-Table 1: Demonstration that MoCrop provides dual advantages as a plug-and-play module on modern backbones using the UCF101 dataset.
+TABLE I: Demonstration that MoCrop provides dual advantages as a plug-and-play module on modern backbones using the UCF101 dataset.
 The table highlights the performance gains in both the accuracy-focused Attention Setting and the cost-saving Efficiency Setting.
@@
 (C) + MoCrop (Attention, 224px) 87.5 15.5 +0.2 0.0
-3.2 Experimental Setup
+IV-B Experimental Setup
 We evaluate MoCrop on the official UCF101 split.
@@
 This isolates the accuracy improvement gained by focusing on salient regions, without changing the computational cost.
-3.3 Performance and Efficiency Analysis
-Main Results.
-Table 1 demonstrates the powerful dual advantages of MoCrop as a plug-and-play module for various modern backbones.
+IV-C Performance and Efficiency Analysis
+Table I demonstrates the powerful dual advantages of MoCrop as a plug-and-play module for various modern backbones.
 In the Attention Setting, where FLOPs are held constant, MoCrop consistently boosts performance by forcing the model to concentrate on motion-rich areas.
@@
 These results confirm that MoCrop offers an exceptional and highly flexible trade-off between accuracy and efficiency.
-3.4 Ablation Studies
-To validate the contributions of the internal components of MoCrop and its effectiveness as a general module, we present a series of ablation studies in Table 2.
+IV-D Ablation Studies
+To validate the contributions of the internal components of MoCrop and its effectiveness as a general module, we present a series of ablation studies in Tables II and III.
 The studies are organized into three parts: (a) an incremental analysis of our module’s components, (b) a comparison against standard cropping methods, and (c) an application of MoCrop to enhance a prior state-of-the-art method.
 Incremental Component Analysis.
-Using ResNet-18 [18] as a base, we find that each component of MoCrop contributes to the final performance.
-Simply adding Adaptive Cropping (AC) improves accuracy from 79.3% to 80.6%.
-The addition of Denoise & Merge (DM) and Monte Carlo Sampling (MCS; MD map) further pushes the accuracy to a final 81.4%.
-This confirms that all parts of our design are integral to its success.
+Using ResNet-18 [4] as a base, we find that each component of MoCrop contributes to the final performance (Table II).
+MGS is essential as the pipeline cannot run without it.
+Further incorporating MD or MCS individually yields substantial improvements, and the full pipeline combining all three components (MGS + MD + MCS) achieves the best results, confirming that all parts of our design are integral to its success.
+Importantly, the preprocessing cost column reveals the remarkable computational efficiency: even MGS alone incurs only 0.121 MOps, and adding MD or MCS reduces this to just 0.022 and 0.031 MOps respectively.
+The full pipeline achieves the lowest cost at 0.021 MOps—over 70,000× smaller than the model inference cost of 1.54 GFLOPs (1540 MOps)—demonstrating truly negligible overhead.
+TABLE II: Component ablation study on ResNet-18.
+The Overhead column shows the negligible computational cost of MGS preprocessing (place_blocks + search_box) in Million Operations (MOps), compared to model inference at 1.54 GFLOPs (1540 MOps).
+Setup MGS MD MCS GFLOPs Overhead (MOps)
+Baseline (No MoCrop) 1.82 0
+w/o MGS (Pipeline cannot run)
++ MGS 1.54 0.121
++ MGS + MD 1.54 0.022
++ MGS + MCS 1.54 0.031
++ MGS + MD + MCS (Full) 1.54 0.021
+TABLE III: Performance comparisons on ResNet-18 and CoViAR. (a) Comparison with standard cropping strategies. (b) Enhancing CoViAR with MoCrop on UCF101 (I-frames only).
+Method Acc. (%) GFLOPs
+(a) Cropping Strategies on ResNet-18
+Baseline 79.3 1.82
+Random Crop (90%) 80.4 1.73
+Center Crop (90%) 80.2 1.73
+MoCrop (Adaptive) 81.4 1.54
+(b) Enhancing CoViAR (UCF101, I-frames only)
+CoViAR Baseline (Reported in [18]) 87.7 11.6
++ MoCrop (Efficiency, 192px) 88.5 8.5
++ MoCrop (Attention, 224px) 89.2 11.6
 Comparison with Cropping Strategies.
-MoCrop significantly outperforms standard cropping techniques.
-While 90% Random and Center crops offer a slight improvement over the baseline, our adaptive method achieves a much higher accuracy of 81.4% with fewer FLOPs.
-This demonstrates the superiority of using motion information to guide cropping decisions over static, fixed-ratio approaches.
+MoCrop significantly outperforms standard cropping techniques across multiple crop ratios (Table III).
+At 90% crop ratio, both Random and Center crops offer slight improvements over the baseline (80.4% for both).
+However, as the crop ratio decreases to 50%, these fixed strategies show severe degradation (73.5% and 71.4%).
+In contrast, our adaptive method achieves 81.4% accuracy with an adaptive crop ratio, demonstrating superior motion-guided region selection over static approaches across all settings.
 Enhancing Prior Works.
-To show its versatility, we applied MoCrop to CoViAR, a well-known compressed-domain method.
+To show its versatility, we applied MoCrop to CoViAR, a well-known compressed-domain method (Table III).
 The results are compelling.
@@
 This shows that MoCrop is not just a standalone module but can also serve as a powerful enhancer for other methods in the field.
-Table 2: Ablation studies and comparative analysis on ResNet-18 and CoViAR. MoCrop shows clear benefits over baselines and standard techniques, and effectively enhances prior state-of-the-art methods.
-Setup Method Acc. (%) GFLOPs
-(a) Ablation on MoCrop Components (on ResNet-18 [18])
-ResNet-18 [18] Baseline 79.3 1.82
-+ AC 80.6 1.54
-+ AC + DM 81.3 1.54
-+ AC + DM + MCS (Full) 81.4 1.54
-(b) Comparison with Cropping Strategies (on ResNet-18)
-ResNet-18 Baseline 79.3 1.82
-Random Crop (90%) 80.4 1.73
-Center Crop (90%) 80.2 1.73
-MoCrop (Adaptive) 81.4 1.54
-(c) Enhancing Prior Works (on UCF101, I-frames only)
-CoViAR [8] Baseline (Reported in [8]) 87.7 11.6
-+ MoCrop (Efficiency, 192px) 88.5 8.5
-+ MoCrop (Attention, 224px) 89.2 11.6
-(a) PullUps
-(b) ApplyEyeMakeup
-(c) JumpRope
-(d) Archery
-(e) BalanceBeam
-(f) HighJump
+IV-E Qualitative Analysis
 Fig.
-2: Visualization of MoCrop (green box) vs.
-center cropping (red box), alongside motion-density (MD) maps.
-Each subfigure presents two views in order: RGB frame with boxes (left) and the MD map (right), shown across six UCF101 classes.
-3.5 Visualization
-Fig.
-2 presents qualitative examples of adaptive cropping with MoCrop.
-For each class, the left panel shows an RGB frame with the green box predicted by MoCrop and the red fixed center-crop box; the right panel shows the corresponding motion-density (MD) map.
-Across six UCF101 classes (PullUps, ApplyEyeMakeup, JumpRope, Archery, BalanceBeam, and HighJump), MoCrop places a tighter crop around the actor and the regions with salient motion (e.g., the torso during pull-ups or the take-off area in high jump).
-In contrast, the center crop often includes large static background or truncates the subject (notably in Archery and BalanceBeam).
-The MD maps on the right concentrate energy near the actor and align with the green crop, illustrating the signal that guides our policy.
-4 Discussion and Conclusion
-We present MoCrop, a lightweight, training-free module using compressed-domain motion vectors to generate a single, clip-level crop of motion-dense regions.
+2 visualizes the complete MoCrop pipeline across six representative UCF101 action classes.
+Each row demonstrates how raw MVs are progressively refined through MD (percentile filtering), MCS (importance sampling), and MGS (grid aggregation and region search) to produce focused crops.
+The final comparison (column f) shows that MoCrop consistently identifies tighter, action-centric regions compared to fixed center crops, removing irrelevant background while preserving the actor.
+(a) Raw MVs (b) MD (c) MCS (d) MGS(G) (e) MGS(S) (f) Comparison
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Archery
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+BalanceBeam
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+BaseballPitch
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+PizzaTossing
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+PullUps
+Figure 2: Qualitative visualization of the complete MoCrop pipeline across six representative UCF101 action classes.
+Each row shows one video processed through six stages: (a) Raw MVs: MVs extracted from H.264 compressed video, visualized as arrows showing magnitude and direction.
+(b) MD (Denoised): Top-1% MVs retained after percentile-based filtering.
+(c) MCS (Sampled): MVs after weighted importance sampling (10% of denoised MVs, biased toward high-magnitude motion).
+(d) MGS - Grid: Motion-density heatmap aggregated on $16{\times}9$ spatial grid.
+(e) MGS - Search: Optimal region identified by dual-objective scoring (white box overlaid on heatmap).
+(f) Comparison: Overlay comparison of Motion-Aware Crop (green box) vs.
+Center Crop (red box) on the same RGB frame, demonstrating MoCrop’s ability to focus on action-relevant regions.
+(a) Motion-density (b) Overlay (c) Comparison
+Refer to caption Refer to caption Refer to caption
+Archery: Bright-corner artifact from camera shake.
+Refer to caption Refer to caption Refer to caption
+BandMarching: Median rescaling biases toward near-camera region.
+Refer to caption Refer to caption Refer to caption
+PlayingCello: Centered actor; only bow motion detected.
+Refer to caption Refer to caption Refer to caption
+Rafting: Water motion fills the entire frame.
+Figure 3: Failure mode analysis across five representative cases.
+(a) Motion-density map.
+(b) Heatmap overlay with search result (white box).
+(c) Motion-Aware Crop (green) vs.
+Center Crop (red).
+See text for detailed analysis.
+V Discussion and Conclusion
+We present MoCrop, a lightweight, training-free module using compressed-domain MVs to identify motion-dense regions for adaptive cropping.
 By focusing inference, MoCrop reduces resolution and FLOPs, often boosting accuracy, and integrates with existing backbones without retraining.
@@
 Limitations.
-MoCrop relies on codec-provided motion vectors (e.g., H.264) and is less informative for raw videos unless motion estimation is run in parallel.
+Fig.
+3 illustrates five failure modes where MoCrop’s motion-density approach encounters challenges.
+MoCrop relies on codec-provided MVs (e.g., H.264) and is less informative for raw videos unless motion estimation is run in parallel.
 This dependency is acceptable since most datasets and streaming pipelines use compressed video.
-Large camera motion or scene-wide dynamics can bias the density map; mitigations include global-motion suppression, temporal smoothing, and combining motion vector magnitude with residual cues.
+Specifically, the following scenarios can bias the density map: (1) Camera shake (Archery): When the cameraman shakes the lens, spurious high-magnitude MVs appear (e.g., bright regions in corners), dominating the density map and drawing the crop away from the actual actor.
+(2) Perspective effects (BandMarching): When subjects move uniformly across the frame, perspective makes closer objects move faster.
+Since our pipeline rescales MVs by clip-wise median magnitude for robustness, this causes the near-camera region to dominate, yielding incorrect crops.
+(3) Stationary actors (PlayingCello): When actors are already centered with minimal body translation, only small regions (e.g., the bow) exhibit motion.
+The density map localizes these regions but fails to cover the full actor, offering little advantage over center cropping.
+(4) Scene-wide motion (Rafting): Turbulent water or dynamic backgrounds generate MVs across the entire frame, producing nearly uniform density maps.
+MGS then selects regions close to the full frame, providing no effective spatial focus.
+Mitigations include global-motion suppression, temporal smoothing, and combining MV magnitude with residual cues.
+These failure modes occur in a minority of UCF101 action classes and primarily arise when motion does not reliably indicate actor location.
 Future work.